Topic: Differentiation.
Question: The average cost function is given by: C(q) = 1 +(2q-4)^2
(a) Calculate the Marginal Cost at q=q0
(b) Calculate q0
Note 0= Zero
Thanks.

Answer by stanbon(75887)

Differentiation.
Question: The average cost function is given by: C(q) = 1 +(2q-4)^2
(a) Calculate the Marginal Cost at q=q0
---
C'(q) = 2(2q-4)*2 = 4(2q-4) = 4q-8
C'(qo) = 4qo-8
----------------------
(b) Calculate qo
Not sure of the conditions for finding qo.
Cheers,
Stan H.